ment Scoring time left 1:52: 15 of 1 100 N (about 248 lb). She pushes the weight to her ful and comes to rest. (b) Assume she is able to move the push-plate of tne leg-press machine at an acceleration of o.s0 m/s2 for the first half of the displacement. What is the normal force on each foot during this period of acceleration? Assume the resistive force is due to a mass with weight of 1100 N hanging on the cable of the leg-press apparatus.
Answer #1

a) Force on each leg= 1100/2= 550 N

b) normal force on each leg= 0.5*m(g-a)= 0.5* 112.49*(9.8+ 0.5)= 579.323 N

Plz rate